OpenStudy (southern_princess13):

MEDAL AND FAN FOR ANSWER!! What is the approximate distance around the half circle with radius 35 inches? A. 1923.25 inches B. 219.8 inches C. 109.9 inches D. 54.95 inches

OpenStudy (legomyego180):

pi*35=?

OpenStudy (legomyego180):

circumference of a circle:\[ = 2\pi r\] Circumference of half a circle = \[\frac{ 1 }{ 2 }(2 \pi r)\] Threrfore your answer is pi times 35

OpenStudy (southern_princess13):

sorry i was doing somthing and thank you :) @legomyego180